OpenStudy (anonymous):

the length of the sides of a triangle are represented by three consecutive odd integers. If the perimeter of the triangle is 75 feet, find the lengths of its sides

OpenStudy (anonymous):

X X+2 X+4 3 SIDES X + X + 2 +X +4 = 75 3X = 75 - 6 3X = 69 X = 23 X+2 = 25 X+ 4 = 27 23 25 27 AER THE THREE SIDES